The riverboat Hiawatha goes 30 km upstream and back in 4 hours. If the Hiawatha's speed in still water is 16 km'h, what is the speed of the current?
Answer by scott8148(6628)

t = d / r ___ total time is 4 h

[30 / (16 - r)] + [30 / (16 + r)] = 4

480 + 30r + 480 - 30r = 1024 - 4r^2

r^2 - 16 = 0